OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos x

OpenStudy (austinl):

sin2x = 2sinxcosx

OpenStudy (anonymous):

I'm still extremely lost....

OpenStudy (austinl):

\[\sin(2x) = 2\sin(x)\cos(x)\] \[\sin(2x) - \cos(x)\rightarrow 2\sin(x)\cos(x) - \cos(x)\]

OpenStudy (anonymous):

2 sin x cos^2x ?

OpenStudy (austinl):

That is a minus sign, that should be your final answer. That I gave.